OpenStudy (anonymous):

A right triangle has one side of length 11 and a hypotenuse of length 12. How long is the third side of the triangle?

OpenStudy (anonymous):

Let lenth of other leg be "a". Therefore, a^2+12^2=20^2 a^2=400-144=256 Therefore, a^2=256 a=16

OpenStudy (anonymous):

woahh im so lost

OpenStudy (anonymous):

Do you know / understand Pythagoras' theorem?

OpenStudy (anonymous):

nope i have failed geometry 3 times

OpenStudy (anonymous):

it's very important to learn and understand as it underpins a lot of geometry at your level. It says in a right angled triangle, if you square the two sides adjacent to the right angle, their sum is equal to the square of the hypotenuse (the side opposite the right angle) It's hard to describe without a picture, it would be worth googling if you dont understand my explanation.

OpenStudy (anonymous):

http://www.mathsisfun.com/pythagoras.html go to this site it makes it easy to understand and kinda fun